The base of a rectangle is twice the side of a square, and the height of the rectangle is two more than the side of the square. The area of the rectangle is 32 square inch units more than the area of the square. Find the dimensions of the rectangle.
Answer by checkley75(3666)

THE SQUARE AREA =X^2
THE AREA OF THE RECTANGLE = 2X(X+2)
2X(X+2)-32=X^2
2X^2+4X-32=X^2
2X^2-X^2+4X-32=0
X^2+4X-32=0
(X-4)(X+8)=0
X-4=0
X=4 ANSWER FOR THE SIDES OF THE SQUARE.
SIDES OF THE RECTANGLE ARE 2*4=8 & 2(4+2)=2*8=16. THE SIDES OF THE RECTANGLE ARE 8 & 16.
PROOF
2*4(4+2)-32=4^4
8*6-32=16
48-32=16
16=16